ARL Weekly News – September 8, 2017


The NOAA Technical Memorandum “Project Sagebrush Phase 2” by Dennis Finn et al. has completed the review and editing process and will shortly be assigned a report number and Digital Object Identifier (DOI) for final publication. This report is a detailed description of the meteorological and tracer observations collected during the light-wind experiments conducted in 2016 at the Idaho National Laboratory. A journal manuscript that references the technical memorandum will be submitted for review shortly.

Completion of the ERT contract renewal at the beginning of September will allow the company to move ahead with filling a vacant scientist position at FRD. This position will be supporting the division’s dispersion activities and will also be assisting with boundary layer research.

Two visitors from the Oak Ridge National Laboratory (ORNL) in Tennessee, including former ATDD employee Kevin Birdwell, will be visiting FRD next week to discuss the HYRad modeling system based on HYSPLIT. HYRad is a candidate system being considered for radiological dispersion modeling applications at ORNL. One of the HYRad updates that will be discussed during the visit is a new user interface Brad Reese has been working on. It is based on Javascript libraries (e.g., Webix and jQuery) rather than the Adobe Flash used in the existing HYRad interface. Flash provided by far the best web interface performance when HYRad was first developed but is now being rapidly depreciated on common browsers.


The wildland fire that started last week continued to slowly burn through the weekend.  The fire would diminish/laydown during the evening and then pop up in the afternoon.  By Monday afternoon, the fire department declared the fire out.  James Wood continued to provide a morning weather forecast with an afternoon update over the weekend. A low level weather surveillance continued over the weekend from home.

Another fire started from a lighting strike this week Wednesday, but was quickly extinguished.

SORD Positions:  Several candidates for the Electronics Technician position were interviewed by phone this week.  The information is being reviewed with a decision to be made next week.  We received the panel for the Meteorologist III position, and will begin reviewing the candidates.  The Job Announcement for the Meteorologist III/IV closed this week.  We look forward to receiving the panel in a couple of weeks and reviewing the candidates.

3D Sonic Anemometers: One of the two sonic anemometers was reset and put back into operation this week.  The sonic lost its programming and was not providing data.  The other anemometer was checked out but was not able to be reset and will need to be replaced next week and returned to RM Young for repair.